OpenStudy (anonymous):

Can this be simplified? cos(180° + x) + cos(180° - x)

OpenStudy (princeharryyy):

yeah !!!

OpenStudy (princeharryyy):

2cos((180+x+180-x)/2)cos((180+x-180+x)/2) = 2*(-1)*cosx = -2cosx

OpenStudy (princeharryyy):

@Katieholmes

myininaya (myininaya):

hint: use sum and difference identity for cosine

myininaya (myininaya):

well that hint leads to another way to go about this i'm trying to determine what identity you used @princeharryyy i'm not saying it is wrong i just want to know the identity and also it might be important to the asker what identity you used

OpenStudy (princeharryyy):

cosA+cosB = 2cos((A+B)/2)cos((A-B)/2)

OpenStudy (princeharryyy):

@myininaya

OpenStudy (princeharryyy):

there's another soln as well but this one is the most easiest one :)

myininaya (myininaya):

well the sum and difference identities are pretty easy to use i haven't used the identities you mentioned in awhile good strategy though

OpenStudy (princeharryyy):

u can also use cos(A+B) and cos(A-B) identities to solve the problem @myininaya

OpenStudy (anonymous):

can you tell me out of this chart which im using?

OpenStudy (princeharryyy):

wat do you want to know in this

OpenStudy (anonymous):

which identity im using to solve this equation

OpenStudy (anonymous):

ive never seen one like this so im a bit confused

OpenStudy (princeharryyy):

how would i know which one u are using :D trignometry is too easy if u know as much of them as u would

OpenStudy (princeharryyy):

if u want u can have link of the properties @Katieholmes

OpenStudy (anonymous):

yes please :)

OpenStudy (princeharryyy):

wait a sec

OpenStudy (princeharryyy):

u can have this for ur reference @Katieholmes ... thogh there are many others but i think this would be sufficient for u

OpenStudy (princeharryyy):

???? hey dere ???

OpenStudy (anonymous):

so it would be the second product identity?

OpenStudy (princeharryyy):

sum to product one in the last :)

OpenStudy (anonymous):

okay so cos+cos

OpenStudy (princeharryyy):

the earlie one that u said is just the extension of the last one they are one of the same if u look closely to them u don't need to remember both of them they are actually same see closely :p

OpenStudy (princeharryyy):

earlier*

OpenStudy (anonymous):

okay thank you!

OpenStudy (princeharryyy):

try and feel maths is pretty easy just improve your speed... u r better than others .... nd hardworking ofcourse

OpenStudy (princeharryyy):

@Katieholmes

OpenStudy (anonymous):

haha thank you so the final answer would be -2cosx

OpenStudy (princeharryyy):

yeah and now i am feeling sleepy ... it's 2:15 am so talk u later bieeee @Katieholmes

OpenStudy (anonymous):

bye haha and thank you

OpenStudy (princeharryyy):

:p

OpenStudy (princeharryyy):

welcome